OpenStudy (anonymous):

8x+y=-16 -3x+y=15 Solve by Elimination!!!!

OpenStudy (rulnick):

-11x=31 x=-31/11

OpenStudy (rulnick):

(First line comes from subtracting the first equation from the second equation.)

OpenStudy (anonymous):

Subtract the 2nd equation from the first to eliminate y:\[11x = -31\]\[x = -\frac{31}{11}\]Then plug in this value for x and solve for y:\[8(-\frac{31}{11}) + y = -16\]\[y = -16 + \frac{248}{11}\]\[y = \frac{72}{11}\]